QUESTION 9 Encrypt the message WALDEN using a shift cipher with f(p (p+17) mod 26
Answer #1

shift cipher is one the simplesgt cipher

here each alphabet are assigned with the numerical value starting from a wth 0 anb b with 2 ............ z= 25.

so

f(p) =(p+17) mod 26means each numbers should be icremented by 17 and mod 26 is caluculated

w= 22 ,

f(p)=22+17 =39 mod 26= 13= n

similarly we get

NRCUVE as encrpted message